OpenStudy (sleepyjess):

How many moles of glucose (C6H12O6) are in 5.0 liters of a 2.5 M C6H12O6 solution?

OpenStudy (sleepyjess):

@ganeshie8

OpenStudy (sleepyjess):

@hartnn

OpenStudy (abmon98):

Number of moles=Concentration(mol/dm^3)*Volume(dm^3)

OpenStudy (anonymous):

M=mole of the solute/Liter of the solution

OpenStudy (abb0t):

Just to clarify on what they have tried to explain here is that you can use what they call "dimensional analysis" here to eliminate some things. Do you remember algebra, and multiplication of them? For example, if you have: \(\sf \frac{a}{b} \times \frac{b}{c} = \frac{a}{c} \) because you notice that the two \(\sf \color{red}{b}\)'s cancel. Well, it's the same here. You're given volume and Molarity, which is \(\sf \large \color{blue}{\frac{mol}{L}}\) So, you multiply by volume, to get moles. Try it, you will see that the units, Liters, cancels out nicely.